On the basis of current legislation, if you earned an average salary of $25,000 over your working life and you retire after the year 2005 at age 62,63, or 64, then your annual Social Security Benefit will be $7000, $7500, or $8000, respectively (Social Security Administration, www.ssa.gov). Write a formula that gives annual benefit as a linear function of age of these three ages.

Answer by stanbon(48568)

You have three pairs of numbers. You only need two to
determine a linear equation.
Points are (62,7000), (63,7500), (64,8000)
Using the 1st two you have a slope of [500/1]
Using x=63, y=7000, m=500 use y=mx+b to find "b".
7000=500(63)+b
7000=31500+b
b=-24500
Your equation is y=500x-24500
Cheers,
Stan H.